I'm having a hard time figuring out this literal equation 2y-2/3x+6=0 for y.
How do you get an answer for y?

Answer by gonzo(654)

2y - 2/3x + 6 = 0
add 2/3x to, and subtract 6 from, both sides of equation:
2y = 2/3x - 6
divide both sides of equation by 2:
y = 2/6x - 3
-----